OpenStudy (anonymous):

What is the product of the polynomials below? (x - 7)(x2 + 3x - 3)

OpenStudy (callisto):

(x - 7)(x^2 + 3x - 3) = x (x^2 + 3x - 3) - 7(x^2 + 3x - 3) = ?

OpenStudy (anonymous):

x3 - 4x2 - 3x + 21

OpenStudy (callisto):

The third term is not correct?!

OpenStudy (anonymous):

x3 - 4x2 - 24x + 21

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

can you help me with my next